# Geometry-Aware Preference Learning for 3D Texture Generation

### v1

@@ Abstract
Recent advances in 3D generative models have achieved impressive results but 3D contents generated by these models may not align with subjective human preferences or task-specific criteria. Moreover, a core challenge in the 3D texture generation domain remains: most existing approaches rely on repeated calls to 2D text-to-image generative models, which lack an inherent understanding of the 3D structure of the input 3D mesh object. To address this, we propose an end-to-end differentiable preference learning framework that back-propagates human preferences, represented by differentiable reward functions, through the entire 3D generative pipeline, making the process inherently geometry-aware. We demonstrate the effectiveness of our framework using four proposed novel geometry-aware reward functions, offering a more controllable and interpretable pathway for high-quality 3D content creation from natural language.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: An overview of the proposed training process, consisting of two main stages: (i) texture generation (Section 2.1), where a latent diffusion model generates high-quality images from textual prompts. Combined with differentiable rendering and 3D vision techniques, this step produces realistic textures for 3D objects. (ii) texture preference learning (Section 2.2), where an end-to-end differentiable pipeline fine-tunes the pre-trained diffusion model $\epsilon_{\theta}$ by maximizing a differentiable reward function $r$. Gradients are backpropagated through the entire 3D generative pipeline, making the process inherently geometry-aware. To demonstrate the method’s effectiveness in producing textures aligned with 3D geometry, we introduce four novel geometry-aware reward functions, detailed in Section 2.3.
@@
While large-scale generative computer vision models learn broad knowledge and some reasoning skills to generate images (Black et al., 2024; Fan et al., 2023), videos (Wu et al., 2023), and 3D objects (Poole et al., 2022; Lin et al., 2023), achieving precise control of their behavior is difficult due to the completely unsupervised nature of their training. Among these generative models, 3D generative ones often rely on adapting text-to-image models at inference (Poole et al., 2022; Lin et al., 2023; Zeng et al., 2024), but the resulting 3D content may not align with human preferences or task-specific needs. This challenge highlights the need for a preference learning frameworks adapted to 3D content creation. One general solution proposed mainly in the large language models’ literature (Ouyang et al., 2022; Christiano et al., 2017; Dong et al., 2023; Rafailov et al., 2023) is to incorporate human or automated feedback as a supervisory signal to guide the generative models toward desired behavior. This is usually done by leveraging reinforcement learning from human feedback (RLHF) (Christiano et al., 2017). Prior work has shown that RLHF can significantly improve results in some domains, such as text generation (Ouyang et al., 2022; Dong et al., 2023) and image synthesis (Black et al., 2024; Fan et al., 2023). However, to the best of our knowledge, DreamReward (Ye et al., 2024) is the only prior work that incorporates human feedback into 3D generative model training. It applies differentiable rewards to optimize NeRF-based geometry editing, without training the generative model. In contrast, we directly fine-tune the diffusion model for texture editing within a 3D generation pipeline, enabling precise control aligned with user preferences. Moreover, unlike DreamReward’s aesthetic rewards that ignore geometry, our geometry-aware rewards ensure textures are both perceptually rich and structurally aligned with the geometry of the input 3D mesh. Despite recent advances, a core challenge in 3D texture generation remains: most existing approaches rely on repeated calls to 2D text-to-image models, which lack an inherent understanding of 3D structure. Therefore, applying preference learning directly to these 2D components often leads to solutions that neglect crucial 3D constraints. To address this, we propose an end-to-end differentiable preference learning framework that back-propagates through the entire 3D generative pipeline, making the process inherently geometry-aware. Our method is reinforcement learning–free, easy to implement, and computationally efficient. By directly fine-tuning the 3D texture generation model through differentiable geometry-aware rewards, our framework supports interactive and user-aligned texture editing, producing results that are both visually compelling and geometrically coherent. By proposing such an approach, we offer a main advantage over existing methods in other generative domains: it is computationally more efficient as it does not necessitate too many expensive sampling steps from the models, which is typically required in reinforcement learning pipelines. The main contributions of this study are: (i) design and development of the first preference learning framework for high-quality 3D geometry-aware content (texture) creation from text, and (ii) proposing novel geometry-aware reward functions for aligning 3D generation models.
@@ 2 Methodology
We propose a simple, reinforcement-learning-free, end-to-end differentiable preference learning framework for generating texture images aligned with the geometric features of a 3D mesh. The training consists of two stages: (i) texture generation, where a diffusion model (Rombach et al., 2022) combined with 3D vision techniques like differentiable rendering (Laine et al., 2020) produces realistic textures from text prompts (Section 2.1); and (ii) texture preference learning, where the diffusion model $\epsilon_{\theta}$ is fine-tuned via a differentiable pipeline that maximizes a reward function $r$, enabling gradients to flow through the full 3D generation process and making it geometry-aware (Section 2.2). We introduce four novel geometry-aware reward functions to demonstrate the method’s effectiveness, each detailed in Section 2.3. An overview of the training procedure of the the proposed end-to-end differentiable framework is shown in Figure 1 and Algorithm 1. The framework supports any form of human preference that can be formulated as a differentiable feedback signal either in the 2D texture space or directly on the 3D surface. This flexibility makes the approach broadly applicable and extensible to a wide range of 3D tasks beyond texture generation.
@@ 2.1 Texture Generation
Our depth-aware text-to-texture pipeline builds on two established methods, (Richardson et al., 2023; Tang et al., 2024). Using a differentiable renderer (Laine et al., 2020), we render the object from multiple viewpoints to extract inpainted images, depth maps, normal maps, and UV coordinates. Given a text prompt, each view is painted using a depth-aware text-to-image diffusion model (Rombach et al., 2022), guided by a pre-trained ControlNet (Zhang et al., 2023) that provides depth information [^1] Pre-Trained ControlNet model. This ensures the generated textures align with both the text and depth (geometry) information. We repeat this process iteratively across all viewpoints until the full 3D surface is painted. Due to the stochastic nature of diffusion models, this can result in inconsistent textures and visible seams. To address this, we adopt a dynamic partitioning strategy inspired by (Richardson et al., 2023; Chen et al., 2023), using the view direction cosine (see Appendix A) to segment each view into three regions: (i) generate – newly visible and unpainted areas, (ii) refine – previously painted areas now seen from a better angle (higher cosine), and (iii) keep – already well-painted regions that require no update. This partitioning guides the rendering and repainting process to maintain visual consistency. However, this partitioning and certain other operations in the texture generation step are not inherently differentiable, so we introduce mathematical modifications to ensure differentiability. These modifications are essential not only for making the pipeline differentiable, enabling the training of the diffusion model, but also for incorporating gradient information from each component of the texture-generation process. This ultimately makes the entire pipeline aware of the geometry of the mesh being textured. For more on the differentiable mathematical formulation, we refer the reader to Section A.1 and Section A.2.
2.2 Texture Preference Learning
The texture image, obtained from the texture generation step, serves as an input to the next step, texture preference learning, to be evaluated and then enhanced according to a differentiable reward which represents human preferences or task-specific objectives. The goal is to fine-tune the pre-trained stable diffusion model’s parameters $\theta$ such that the texture images maximize a differentiable reward function $r$. The maximization problem can be then formulated as:
@@ 2.2 Texture Preference Learning
where $r$ is a differentiable reward function, $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}})$ represents the iterative texturing process from different viewpoints $v_{gen}$ (explained in Section 2.1 and Appendix A), and $c$ is the text prompt randomly sampled from the text dataset $p_{c}$. To solve Equation 1, instead of maximizing the reward function, we consider the negative of it as a loss function $L(\theta)=-J(\theta)$. Then, we compute $-\nabla_{\boldsymbol{\theta}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}}\right),\mathbf{c}\right)$ and update $\theta$ using gradient-based optimization algorithms and the following update rule: $\theta\leftarrow\theta+\eta\nabla_{\boldsymbol{\theta}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}}\right),\mathbf{c}\right)$. However, naively optimizing the $\theta$ would results in requiring a massive amount of memory which makes the fine-tuning process impractical. This is because of the fact that computing the gradients $\nabla_{\boldsymbol{\theta}}r$ requires backpropagation through multiple diffusion steps and multiple camera viewpoints which stems from the texturing process $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}})$. Consequently, this necessitates storing all the intermediate variables during the multi-view texturing process. More specifically, each step in the texturing process involves at least 10 diffusion steps, and with 8–10 total texturing steps (corresponding to the number of camera viewpoints), this results in approximately 100 diffusion steps overall. Each step requires around 3 GB of GPU memory, leading to a total memory demand of roughly 300 GB to train the texture-generation pipeline. Such hardware requirements are nearly infeasible, especially in typical academic environments. To alleviate this issue, we leverage two main approaches to reduce the memory used by our method during the training stage: 1) low-rank adaptation (LoRA) (Hu et al., 2022) and 2) gradient checkpointing (Chen et al., 2016). Applying LoRA to the U-Net in the latent diffusion model reduces trainable parameters by roughly a factor of 1000 compared to full fine-tuning. Additionally, two levels of activation checkpointing keep memory usage constant across views and diffusion steps, enabling higher-resolution renders and more diffusion steps. Further memory-saving details are provided in Appendix B.
@@ 2.3 Geometry-Aware Reward Design
Our differentiable framework enables the design and implementation of task-specific differentiable reward functions (which represent human preferences), defined either in the 2D texture image space or directly on the 3D mesh surface, to guide the texture generation toward the desired outcome. Hence, we propose four novel geometry-aware reward functions to demonstrate the effectiveness of the proposed differentiable framework, each detailed in the following subsections.
Algorithm 1 Differentiable 3D Texture Preference Learning
@@
0:  Inference timesteps $T$, training iterations $E$, prompt dataset $p_{c}$, camera viewpoints for texture generation $V_{gen}$, camera viewpoints for texture evaluation $V_{eval}$, pre-trained diffusion model $\epsilon_{\theta}$ conditioned on control commands generated by the ControlNet model, and diffusion model weights $\theta$
  for $e=1$ to $E$ do
     # Step 1: Texture Generation
     Sample a prompt $c\sim p_{c}$, and sample $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
     Initialize the 3D asset’s parameters $a$ including texture image $albedo$, weight image $cnt$, and the view cosine cache image $viewcos$.
     for $v_{gen}$ in $V_{gen}$ do
        $h^{i}=DiffRender(a,v_{gen})$
        for $t=T$ to $1$ do
           $y^{i}_{t-1}=ControlNet(h^{i},t,c)$
           $x^{i}_{t-1}=\mu(x^{i}_{t},y^{i}_{t-1},t,c)+\sigma_{t}z,\;\;\;z\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
        end for
        $Inpaint3DAsset(a,v_{gen})$
        $albedo\leftarrow Update3DAsset(a,v_{gen})$
     end for
     # Step 2: Texture Preference Learning
     if 3D Reward then
        for $v_{eval}$ in $V_{eval}$ do
           $h^{i}=DiffRender(a,v_{eval})$
           $reward=reward+r(h^{i})$
        end for
        $reward=reward/len(V_{eval})$
        $g=-\nabla_{\theta}reward$
     else if Texture Reward then
        $g=-\nabla_{\theta}r(albedo)$
     end if
     $\theta\leftarrow\theta-\eta g$
  end for
  return $\theta$
2.3.1 Geometry-Texture Alignment Reward
@@
To encourage texture image features (e.g. edges) to be aligned with the geometry of a 3D object, we introduce a differentiable reward function that aligns texture gradient vectors with principal curvature directions on the mesh surface (see Appendix C for details). These curvature vectors encode the surface’s bending behavior, and aligning texture gradients with them results in textures that are both visually and semantically consistent with the underlying geomtery of the 3D input mesh. More specifically, for each point in the UV space (texture), obtained by projecting 3D mesh vertices into 2D texture space, we compute the alignment as the squared dot product (cosine similarity) between the normalized texture gradient and corresponding normalized curvature direction vector:
@@
where $N$ is the number of UV coordinates, $\overrightarrow{\mathbf{g}}_{i}$ and $\overrightarrow{\mathbf{c}}_{i}$ are the texture gradient and minimum principal curvature at each UV coordinate, respectively.
Note that curvature is a vertex-level attribute defined per vertex on the 3D mesh surface, whereas texture gradients are defined per pixel in the 2D UV space. Therefore, to guide texture generation using surface-aware vector fields and to compute cosine similarity between curvature vectors and texture gradients (Equation 2), both must be in the same space. To address this, we project the minimum principal curvature direction vectors, defined at each vertex on the 3D mesh, into the corresponding 2D UV space. Unlike scalar quantities like mean curvature, vector fields require special process to preserve both direction and tangency. This process involves the following steps.
Compute principal curvature directions. We use the method from libigl (Jacobson et al., 2018) to compute the minimum principal curvature direction $d_{i}\in\mathbb{R}^{3}$ at each vertex $v_{i}\in\mathbb{R}^{3}$ of the mesh. Each direction vector lies in the tangent plane of the surface at its vertex.
Project to the tangent plane. Although the direction vectors are already tangent to the surface by construction, we explicitly project them onto the local tangent plane to remove residual normal components (due to numerical noise). For each vertex: $\mathbf{d}_{i}^{\tan}=\mathbf{d}_{i}-\left(\mathbf{d}_{i}\cdot\mathbf{n}_{i}\right)\mathbf{n}_{i}$ where $\mathbf{n}_{i}$ is the surface normal at vertex $i$.
Trace vector to find triangle containment. For each projected vector $\mathbf{d}_{i}^{\tan}$, we trace it forward from the verex $v_{i}$ to obtain a new point: $\mathbf{p}_{i}=\mathbf{v}_{i}+\frac{1}{\lambda_{i}}\cdot\mathbf{d}_{i}^{\tan}$ where $\lambda_{i}$ is a normalization constant to ensure the traced vector remains within a local triangle. We then search among the projected triangles incident to $v_{i}$ to find one in which the point $\mathbf{p}_{i}$ lies. If found, we compute the barycentric coordinates $\beta_{i}=[\beta_{1},\beta_{2},\beta_{3}]$ of $\mathbf{p}_{i}$ relative to the triangle’s projected vertices.
Convert to UV coordinates. Using the barycentric weights and the corresponding UV triangle $(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3})$, we map the end of the vector to UV space: $\mathbf{u}_{end}=\beta_{1}\mathbf{u}_{1}+\beta_{2}\mathbf{u}_{2}+\beta_{3}\mathbf{u}_{3}$. The UV origin $\mathbf{u}_{orig}$ is determined from the texture coordinate associated with vertex $v_{i}$. The 2D texture-space direction vector is then:
@@
Populate texture coordinates. Since a single vertex may be associated with multiple UV coordinates (due to UV seams), the final $\mathbf{d}_{i}^{\mathbf{uv}}$ is assigned to all associated texel indices. This step builds a complete UV-space vector field aligned with curvature directions.

### v2

# End-to-End Fine-Tuning of 3D Texture Generation using Differentiable Rewards

@@ Abstract
While recent 3D generative models can produce high-quality texture images, they often fail to capture human preferences or meet task-specific requirements. Moreover, a core challenge in the 3D texture generation domain is that most existing approaches rely on repeated calls to 2D text-to-image generative models, which lack an inherent understanding of the 3D structure of the input 3D mesh object. To alleviate these issues, we propose an end-to-end differentiable, reinforcement-learning-free framework that embeds human feedback, expressed as differentiable reward functions, directly into the 3D texture synthesis pipeline. By back-propagating preference signals through both geometric and appearance modules of the proposed framework, our method generates textures that respect the 3D geometry structure and align with desired criteria. To demonstrate its versatility, we introduce three novel geometry-aware reward functions, which offer a more controllable and interpretable pathway for creating high-quality 3D content from natural language. By conducting qualitative, quantitative, and user-preference evaluations against state-of-the-art methods, we demonstrate that our proposed strategy consistently outperforms existing approaches. We will make our implementation code publicly available upon acceptance of the paper.
{strip}[Uncaptioned image]
Figure 1: Texture images generated before and after reward fine‑tuning on different 3D mesh objects using different rewards. Each column shows the text prompt at the top and the corresponding reward function used for texture generation at the bottom. Texture images produced by our method consistently outperform the pre‑fine‑tuning baseline (InTeX [30]) across all rewards and experiments.
@@ 1 Introduction
Figure 2: An overview of the proposed training process, consisting of two main stages: (i) texture generation (Sec. 3.1), where a latent diffusion model generates high-quality images from textual prompts. Combined with differentiable rendering and 3D vision techniques, this step produces realistic textures for 3D objects. (ii) texture reward learning (Sec. 3.2), where an end-to-end differentiable pipeline fine-tunes the pre-trained text-to-image diffusion model by maximizing a differentiable reward function $r$. Gradients are back-propagated through the entire 3D generative pipeline, making the process inherently geometry-aware. To demonstrate the method’s effectiveness in producing textures aligned with 3D geometry, we introduce five novel geometry-aware reward functions, detailed in Sec. 3.3 and Appendix F.
@@
While large-scale generative computer vision models learn broad knowledge and some reasoning skills to generate images [3, 11], videos [32], and 3D objects [23, 16], achieving precise control of their behavior is difficult due to the completely unsupervised nature of their training. Among these generative models, 3D generative ones often rely on adapting text-to-image models at inference [23, 16, 34], but the resulting 3D content may not align with human preferences or task-specific needs. This challenge highlights the need for a preference learning frameworks adapted to 3D content creation. One general solution proposed mainly in the large language models’ literature [20, 6, 9, 25] is to incorporate human or automated feedback as a supervisory signal to guide the generative models toward desired behavior. This is usually done by leveraging reinforcement learning from human feedback (RLHF) [6]. Prior work has shown that RLHF can significantly improve results in some domains, such as text generation [20, 9] and image synthesis [3, 11]. However, to the best of our knowledge, DreamReward [33] is the only prior work that incorporates human feedback into 3D generative model training. It applies differentiable rewards to optimize NeRF-based geometry editing, without training the generative model. In contrast, we directly fine-tune the diffusion model for texture editing within a 3D generation pipeline, enabling precise control aligned with user preferences. Moreover, unlike DreamReward’s aesthetic rewards that ignore geometry, our geometry-aware rewards ensure textures are both perceptually rich and structurally aligned with the geometry of the input 3D mesh. Despite recent advances, a core challenge in 3D texture generation remains: most existing approaches rely on repeated calls to 2D text-to-image models, which lack an inherent understanding of 3D structure. Therefore, applying preference learning directly to these 2D components often leads to solutions that neglect crucial 3D constraints. To address this, we propose an end-to-end differentiable preference learning framework that back-propagates differentiable rewards through the entire 3D generative pipeline, making the process inherently geometry-aware. By directly fine-tuning the 3D texture generation model through differentiable geometry-aware rewards, our framework supports interactive and user-aligned texture editing, producing results that are both visually compelling and geometrically coherent. The main contributions of this study are:
  • We propose a method to fine-tune a 2D text-to-image foundational model within a 3D generative pipeline in an end-to-end fashion using 3D geometry-aware rewards.
  • We propose several novel geometry-aware rewards that can be used in conjunction with our end-to-end differentiable framework to improve texture generation on challenging 3D shapes and realistic user requirements.Empirically validating their performance.
2 Related Work
Texture Generation.
Modern 3D texture generation methods leverage text-to-image or image-to-image diffusion models [28, 8, 27] to produce high-quality, realistic textures. These techniques broadly fall into two categories:
(i) Texturing via score distillation sampling (SDS). These methods treat 3D shape parameters as learnable and uses SDS as supervision. DreamFusion [23] introduced this by optimizing a NeRF [19] whose views are scored by a pre-trained diffusion model. Latent-NeRF [18] (LatentPaint) extends this by directly optimizing UV texture pixels, and Magic3D [16] accelerates the process via a two-stage coarse-to-fine pipeline. While these methods yield compelling textures, they remain computationally expensive (1–2 hours per asset), often noisy, and offer limited control due to entangled geometry and appearance. (ii) Texturing via direct use of image-generation diffusion models. Alternatively, methods like TEXTure [26], Text2Tex [4], and InTex [30] assume fixed geometry and generate textures by rendering multi-view images, applying a depth-aware diffusion model, and projecting results back to the UV atlas. This decoupling enables faster runtime and easier editing but lacks a deep understanding of 3D structure, often misaligning textures with human intent or task needs. Our work builds on this paradigm by introducing a geometry-aware fine-tuning stage that aligns textures with human preferences and task-specific objectives, improving coherence and control without the overhead of joint optimization.
Human Preference Learning.
Recently popularized in the large language model community [20, 6, 9, 25], this area leverages human or automated feedback as supervision to steer generative models toward desired behaviors. Typically implemented via reinforcement learning from human feedback (RLHF) [6], this approach has significantly advanced text generation [20, 9] and image synthesis [3, 11]. In 3D generation, however, its application remains limited. DreamReward [33] learns NeRF-based shape and appearance using differentiable rewards combined with SDS loss, but due to the entangled nature of NeRF representations, it offers limited control over texture. DeepMesh [36], by contrast, applies direct preference optimization (DPO) [25] to mesh geometry alone, neglecting texture modeling entirely. Our approach departs from both. We focus on explicit mesh texturing, not implicit NeRF reconstruction, and fine-tune the diffusion model within a 3D texture-generation pipeline. By designing geometry-aware reward functions that align texture gradients with surface curvature, our method enables precise, user-driven control over texture appearance. The resulting textures are perceptually rich, structurally coherent, and aligned with human preferences or task-specific objectives via differentiable rewards.
@@ 3 Methodology
We propose a simple, reinforcement-learning-free, end-to-end differentiable framework that can fine-tune a 3D texture generation model to learn any human or task-specific preferences that can be represented as differentiable rewards. The training consists of two stages: (i) texture generation, where a diffusion model [27] combined with 3D vision techniques like differentiable rendering [15] produces realistic textures from text prompts (Sec. 3.1); and (ii) texture reward learning, where the diffusion model is fine-tuned via a differentiable pipeline that maximizes a reward function $r$, enabling gradients to flow through the full 3D generation process and making it geometry-aware (Sec. 3.2). To demonstrate the method’s effectiveness, we introduce five novel geometry-aware differentiable reward functions, each detailed in Sec. 3.3 and Appendix F. An overview of the training procedure of the the proposed end-to-end differentiable framework is shown in Fig. 2 and Algorithm 1. The framework supports any form of human preference that can be formulated as a differentiable feedback signal either in the 2D texture space or directly on the 3D surface. This flexibility makes the approach broadly applicable and extensible to a wide range of 3D tasks beyond texture generation.
@@ 3.1 Texture Generation
Our depth-aware text-to-texture pipeline builds on the methods established in [26] and [30]. Using a differentiable renderer [15], we render the object from multiple viewpoints to extract inpainted images, depth maps, normal maps, and UV coordinates. Given a text prompt, each view is painted using a depth-aware text-to-image diffusion model [27], guided by a pre-trained ControlNet [35] that provides depth information [^1] https://huggingface.co/ashawkey/control_v11e_sd15_depth_aware_inpaint. This ensures the generated textures align with both the text and depth (geometry) information. We repeat this process iteratively across all viewpoints until the full 3D surface is painted. Due to the stochastic nature of diffusion models, this can result in inconsistent textures and visible seams. To address this, we adopt a dynamic partitioning strategy inspired by [26, 4], using the view direction cosine concept (see Appendix A) to segment each view into three regions: (i) generate – newly visible and unpainted areas, (ii) refine – previously painted areas now seen from a better angle (higher cosine), and (iii) keep – already well-painted regions that require no update. This partitioning guides the rendering and repainting process to maintain visual consistency. However, this partitioning and certain other operations in the texture generation step are not inherently differentiable, so we introduce mathematical modifications to ensure differentiability. These modifications are essential not only for making the pipeline differentiable, enabling the training of the diffusion model, but also for incorporating gradient information from each component of the texture-generation process. This ultimately makes the entire pipeline aware of the geometry of the mesh being textured. For more on the differentiable mathematical formulation, we refer the reader to Sec. A.1 and Sec. A.2.
3.2 Texture Rewards Learning
The texture image, obtained from the texture generation step, serves as an input to the next step, texture reward learning, to be evaluated and then enhanced according to a differentiable reward. The goal is to fine-tune the pre-trained stable diffusion model’s parameters $\theta$ such that the texture images maximize a differentiable reward function $r$. The maximization problem can then be formulated as:
@@ 3.2 Texture Rewards Learning
where $r$ is a differentiable reward function, $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}})$ represents the iterative texturing process from different viewpoints $v_{gen}$ (explained in Sec. 3.1 and Appendix A), and $c$ is the text prompt randomly sampled from the text dataset $p_{c}$. To solve Eq. 1, instead of maximizing the reward function, we consider the negative of it as a loss function $L(\theta)=-J(\theta)$. Then, we compute $-\nabla_{\boldsymbol{\theta}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}}\right),\mathbf{c}\right)$ and update $\theta$ using gradient-based optimization algorithms and the following update rule: $\theta\leftarrow\theta+\eta\nabla_{\boldsymbol{\theta}}r\left(\operatorname{TexGen}\left(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}}\right),\mathbf{c}\right)$. However, naively optimizing the $\theta$ would results in requiring a massive amount of memory which makes the fine-tuning process impractical. This is because of the fact that computing the gradients $\nabla_{\boldsymbol{\theta}}r$ requires backpropagation through multiple diffusion steps and multiple camera viewpoints which stems from the texturing process $\operatorname{TexGen}(\boldsymbol{\theta},\mathbf{c},\mathbf{v_{gen}})$. Consequently, this necessitates storing all the intermediate variables during the multi-view texturing process. More specifically, each step in the texturing process involves at least 10 diffusion steps, and with 8–10 total texturing steps (corresponding to the number of camera viewpoints), this results in approximately 100 diffusion steps overall. Each step requires around 3 GB of GPU memory, leading to a total memory demand of roughly 300 GB to train the texture-generation pipeline. Such hardware requirements are nearly infeasible, especially in typical academic environments. To alleviate this issue, we leverage two main approaches to reduce the memory used by our method during the training stage: 1) low-rank adaptation (LoRA) [13], and 2) gradient checkpointing [5]. Applying LoRA to the U-Net in the latent diffusion model reduces trainable parameters by roughly a factor of 1000 compared to full fine-tuning. Additionally, two levels of activation checkpointing keep memory usage constant across views and diffusion steps, enabling higher-resolution renders and more diffusion steps. Further memory-saving details are provided in Appendix B.
@@ 3.3 Geometry-Aware Reward Design
Our differentiable framework enables the design and implementation of task-specific reward functions that encode human preferences, operating in either the 2D texture image domain or directly on the 3D mesh surface, to steer texture generation toward the desired outcome. To illustrate its versatility, we introduce five novel geometry-aware reward functions. First, we present two simple rewards, geometry-aware texture colorization and texture features emphasis, to verify that our pipeline reliably fulfills each reward’s intended objective. We then propose three rewards with potential practical applications, geometry–texture alignment, symmetry-aware texture generation, and camera-pose-aware texture generation, to demonstrate the pipeline’s capacity for producing high-fidelity, realistic, and task-specific textures on 3D shapes. Detailed descriptions of each reward follow in the subsections below and Appendix G. For the geometry-aware colorization reward, we refer the reader to Appendix F.
Algorithm 1 Differentiable 3D Texture Reward Learning
@@
0: Diffusion steps $T$, training iterations $E$, prompt dataset $p_{c}$, viewpoints for generation $V_{gen}$ and evaluation $V_{eval}$, pre-trained model $\epsilon_{\theta}$ and diffusion model weights $\theta$
1:for $e=1$ to $E$ do
2:  # Step 1: Texture Generation
3:  Sample a prompt $c\sim p_{c}$, and sample $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@
4:  Initialize 3D asset parameters $a$: texture $albedo$, weight $cnt$, view cache $viewcos$
5:  for $v_{gen}$ in $V_{gen}$ do
6:   $h^{i}=\mathrm{DiffRender}(a,v_{gen})$
7:   for $t=T$ to 111 do
8:    $y^{i}_{t-1}=\mathrm{ControlNet}(h^{i},t,c)$
9:    $x^{i}_{t-1}=\mu(x^{i}_{t},y^{i}_{t-1},t,c)+\sigma_{t}z,\;\;z\sim\mathcal{N}(\mathbf{0},\mathbf{I})$
10:   end for
[…]
12:   $albedo\leftarrow\mathrm{Update3DAsset}(a,v_{gen})$
13:  end for
14:  # Step 2: Reward Learning
15:  if 3D Reward then
16:   for $v_{eval}$ in $V_{eval}$ do
17:    $h^{i}=\mathrm{DiffRender}(a,v_{eval})$
18:    $reward=reward+r(h^{i})$
19:   end for
20:   $reward=reward/len(V_{eval})$
21:   $g=-\nabla_{\theta}reward$
22:  else if Texture Reward then
23:   $g=-\nabla_{\theta}r(albedo)$
24:  end if
25:  $\theta\leftarrow\theta-\eta g$
26:end for
27:return $\theta$
[…]
where $\nabla\mathrm{T}(x,y)=\sqrt{\left\|\frac{\partial\mathrm{T}}{\partial x}(x,y)\right\|^{2}+\left\|\frac{\partial\mathrm{T}}{\partial y}(x,y)\right\|^{2}}$ is the texture gradient magnitude and $N$ is the total number of pixels in the texture image. Further details about this reward are provided in Appendix E.
@@
Geometry-Texture Alignment Reward. To encourage texture image features (e.g. edges) to be aligned with the geometry of a 3D object, we introduce a differentiable reward function that aligns texture gradient vectors with principal curvature directions on the mesh surface (see Appendix C for details). These curvature vectors encode the surface’s bending behavior, and aligning texture gradients with them results in textures that are both visually and semantically consistent with the underlyinggeomtery of the 3D input mesh. More specifically, for each point in the UV space (texture) obtained by projecting 3D mesh vertices into 2D texture space, we compute the alignment as the squared dot product (cosine similarity) between the normalized texture gradient and corresponding normalized minimum curvature direction vector:
@@
where $N$ is the number of UV coordinates, $\overrightarrow{\mathbf{g}}_{i}$ and $\overrightarrow{\mathbf{c}}_{i}$ are the texture gradient and minimum principal curvature at each UV coordinate, respectively. This alignment works because curvature directions capture the intrinsic anisotropy of the surface, guiding texture flow along perceptually meaningful geometric features such as ridges, valleys, and folds. Note that minimum and maximum curvatures are a vertex-level attribute defined per vertex on the 3D mesh surface, whereas texture gradients are defined per pixel in the 2D UV space. Therefore, to guide texture generation using surface-aware vector fields and to compute cosine similarity between curvature vectors and texture gradients (Eq. 3), both must be in the same space. To address this, we project the minimum principal curvature direction vectors, defined at each vertex on the 3D mesh, into the corresponding 2D UV space. Unlike scalar quantities like mean curvature, vector fields require special process to preserve both direction and tangency. This process involves the following steps.
(I) Computing principal curvature directions. We use the method from libigl [14] to compute the minimum principal curvature direction $\mathbf{d}_{i}\in\mathbb{R}^{3}$ at each vertex $\mathbf{v}_{i}\in\mathbb{R}^{3}$ of the mesh. Each direction vector lies in the tangent plane of the surface at its vertex.
(II) Projecting to the tangent plane. In the next step, the calculation is restricted to the tangent plane of $\mathbf{v}_{i}$ so that we project one-ring adjacent vertices of $\mathbf{v}_{i}$ to the tangent plane. Although the direction vectors are already tangent to the surface by construction, we explicitly project them onto the local tangent plane to remove residual normal components (due to numerical noise). For each vertex: $\mathbf{d}_{i}^{\tan}=\mathbf{d}_{i}-\left(\mathbf{d}_{i}\cdot\mathbf{n}_{i}\right)\mathbf{n}_{i}$ where $\mathbf{n}_{i}$ is the surface normal at vertex $i$.
(III) Tracing vector to find triangle containment. For each projected vector $\mathbf{d}_{i}^{\tan}$, we trace it forward from the vertex $v_{i}$ to obtain a new point: $\mathbf{p}_{i}=\mathbf{v}_{i}+\frac{1}{\lambda_{i}}\cdot\mathbf{d}_{i}^{\tan}$ where $\lambda_{i}$ is a normalization constant to ensure the traced vector remains within a local triangle. We then search among the projected triangles incident to $v_{i}$ to find one in which the point $\mathbf{p}_{i}$ lies. Once found, we compute the barycentric coordinates $\beta_{i}=[\beta_{1},\beta_{2},\beta_{3}]$ of $\mathbf{p}_{i}$ relative to the triangle’s projected vertices.
(IV) Converting to UV coordinates. Using the barycentric weights and the corresponding UV triangle $(\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3})$, we map the end of the vector to UV space: $\mathbf{u}_{\text{end}}=\beta_{1}\mathbf{u}_{1}+\beta_{2}\mathbf{u}_{2}+\beta_{3}\mathbf{u}_{3}$. The UV origin $\mathbf{u}_{\text{orig}}$ is determined from the texture coordinate associated with vertex $\mathbf{v}_{i}$. The 2D texture-space direction vector is then:
@@
(V) Populating texture coordinates. Since a single vertex may be associated with multiple UV coordinates (due to UV seams), the final $\mathbf{d}_{i}^{\mathbf{uv}}$ is assigned to all associated texel indices. This step builds a complete UV-space vector field aligned with curvature directions.